OpenStudy (anonymous):

Which equations are equivalent to the given equation? 49=-7k Choose all answers that are correct. A.49/(-7)=-7k/(-7) B.49*(-7)=-7k/(-7) C.49*(-1/7)=-7k*(-7) D.49*(-1/7)=-7k*(-1/7)

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@NormaV

OpenStudy (jordanloveangel):

ill try no promises

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@jojo4eva

OpenStudy (jordanloveangel):

Solution for 49=-7k equation: Simplifying 49 = -7k Solving 49 = -7k Solving for variable 'k'. Move all terms containing k to the left, all other terms to the right. Add '7k' to each side of the equation. 49 + 7k = -7k + 7k Combine like terms: -7k + 7k = 0 49 + 7k = 0 Add '-49' to each side of the equation. 49 + -49 + 7k = 0 + -49 Combine like terms: 49 + -49 = 0 0 + 7k = 0 + -49 7k = 0 + -49 Combine like terms: 0 + -49 = -49 7k = -49 Divide each side by '7'. k = -7 Simplifying k = -7

OpenStudy (jojo4eva):

A,C, and D

OpenStudy (anonymous):

@jojo4eva can u give @jordanloveangel a medal

OpenStudy (jojo4eva):

uhmm i guess sure

OpenStudy (jordanloveangel):

no need if u dont want 2

OpenStudy (jojo4eva):

Naw i gave you one. You did answer correctly before i did so i gave you one

OpenStudy (jordanloveangel):

thanx

OpenStudy (anonymous):

@jordanloveangel you cheated you went to this site to figure it out and it gives u the answer nice try http://www.geteasysolution.com/